what is the mass (in g) of sodium in a pure sample of 250.0 g of sodium chloride (NaCl) ?
1
Expert's answer
2020-04-18T07:01:31-0400

molar mass of NaCl is 58.5 grams

moles of NaCl = "\frac{250}{58.5}=4.27"

as every molecule of NaCl contains one atom of sodium

so there are 4.27 mole of sodium

mass of 4.27 mole of sodium = 4.27*23 = 98.2 grams